OpenStudy (anonymous):

the design of a digital box camera maximizes the volume while keeping the sum of the dimensions at 6 inches. if the length must be 1.5 times the height, what should each dimension be?

OpenStudy (primeralph):

A+B+C = 6 A*B*C = some number to be maximized A*B = 1.5 From the first equation: Therefore, A+1.5/A +C = 6 C = 6-A-1.5/A From the second equation: A * (1.5/A) * (6-A-1.5/A) = Maximum volume Use derivatives to obtain maxima.